Instrument Log
AeroChem C1 CF Lamont, OK

Date

Technician

Comments

7/14/03

MR

Removed the filters from the system and will send back to the mentor. The system switch will be left in the maintenance position until new filters arrive. Will notify the mentor of the condition.

7/7/03

MR

Did not change the filters due to no filters have arrived from the mentor. Will change as soon as they arrive.

6/30/03

MR

Changed the filters. Status OK.

6/16/03

MR

Changed the filters. Status OK.

6/4/03

MR

Changed the filters. Status OK

5/30/03

MR

Changed the filters. Status OK.

5/26/03

MR

Changed the filters. Status OK.

5/13/03

MR

Changed the filters. Status OK.

5/2/03

MR

Changed the filters in the carousel. Status OK.

4/25/03

MR

Changed the filters. Status OK.

4/14/03

MR

The filters looked as if the system did not sample the filters. Left the filters in the system and will restart the system.

4/7/03

MR

Changed the filters. Status OK.

3/25/03

MR

The system did not recognize the last filter change. Reset the filter change switch.

3/20/03

MR

Changed the filters. Status OK.

3/11/03

MR

The system did not recognize the last filter change. Reset the filter change switch.

3/7/03

MR

The filters looked as if the system did not sample the filters. Left the filters in the system and will restart the system.

2/28/03

MR

Changed the filters. Status OK

2/21/03

MR

The filters looked as if the system did not sample the filters. Left the filters in the system and will restart the system.

2/14/03

MR

Carousel Filters. Status OK

2/4/03

MR

Changed the filters. Status OK

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT. Power was restored to the Central Facility at approximately 1945 GMT. RTS @2000 GMT.

1/21/03

MR

The system did not recognize the last filter change. Reset the filter change switch.

1/17/03

MR

Changed the filters. Status OK

1/10/03

MR

The filters looked as if the system did not sample the filters. Left the filters in the system and will restart the system.

1/3/03

MR

Changed the filters. Status OK

12/27/02

MR

Changed the filters. Status OK

12/20/02

MR

The system did not recognize the last filter change. Reset the filter change switch.

12/16/02

MR

Changed the filters. Status OK

12/2/02

MR

Changed the filters. Status OK.

11/19/02

MR

Changed the filters. Status OK.

11/6/02

MR

Its looks as if the system did not recognize the last filter change and did not sample this filter pack. Will leave the pack in the system and restart the system.

10/30/02

MR

A spare filter pack has arrived on site and needs to be changed out. Changed the filters. Status OK.

10/28/02

MR

The filters need to be changed but there are non available. Will leave the filters in the system until the spare filters arrive.

10/16/02

MR

One spare filter was found in the AOS trailer and needs to be changed out with the one that is in the system. Changed the filters. Status OK.

10/3/02

MR

The filters need to be changed but there are no spare filters. Will wait until spare filters arrive before changing.

9/23/02

MR

Changed the filters. Status OK.

9/11/02

MR

The system email looks as if the system restarted the filter pack. A check of the system found that the system maintenance switch had been flipped by mistake. Reset the switch and checked the filters. The filters looked as if they had not been sampled. Left the filters in the system to be sampled.

9/5/02

MR

The filters looked as if the system did not sample all 7 filters. Changed the filters and will restart the system. Status ok.

8/29/02

MR

Changed the carousel filters. Status ok.

8/26/02

MR

The system e-mail indicates the system did not recognize the last filter change. Flipped the switch to restart the filter sampling.

8/21/02

MR

Changed the filters. Status ok.

8/19/02

MR

The system e-mail indicates that the system is not finished sampling the filters. Will allow the system to finish sampling the rest of the filters before changing.

8/12/02

MR

The system e-mail status looks as if the system did not recognize the last filter change. Flipped the switch to restart the filter sampling.

8/9/02

MR

Changed the carousel filters. Status ok.

8/2/02

MR

The system looked as if the filters where not sampled. Left the filters in the system to be resample.

7/29/02

MR

Changed the carousel filters. Status ok.

7/26/02

MR

The system is due to be changed but according to the system e-mails the filters are still on filter # 6. Will allow the system to run through the rest of the filters before changing them.

7/19/02

MR

The AOS system did not seem to recognize the last filter change. Flipped the maintenance switch to restart the filters.

7/17/02

MR

Received new filters. Changed the filters. Status OK.

7/15/02

MR

No new filters have been returned from the mentor. No filter change. Will notify the mentor that they system needs new filters.

7/8/02

MR

No new filters have been returned from the mentor. No filter change.

7/1/02

MR

No new filters have been returned from the mentor. No filter change.

6/24/02

MR

Changed the carousel filters. Status ok.

6/18/02

MR

According to the system e-mails it looks as if the system did not recognize the filter change on 6/17/02. Flipped the maintenance switch between the maintenance position and run position a couple of times and the returned it to the run position.

6/17/02

MR

Changed the carousel filters. Status ok.

6/13/02

MR

Powered up the AOS system at 1430 GMT.

6/13/02

MR

The power to the site will be shut down for site maintenance. Powered down the AOS system.

6/6/02

MR

Repairs are completed. It looks as if the system has restarted on filter #4. Will allow the system to run until all filters have been sampled.

5/27/02

MR

The AOS system has stopped due to a disk failure. Dick Eagan is on site performing repairs. The system will be down until repairs are completed.

5/24/02

MR

Changed the carousel filters. Status ok.

5/17/02

MR

Changed the carousel filters. Status ok.

5/6/02

MR

The system is now indicating that the last filter has been sampled. Changed the carousel filters. Status ok.

5/2/02

MR

The system e-mails indicate that the system is still trying to sample all the filters and is only on filter #5. Will allow the system to run until the filters are all sampled.

4/25/02

MR

Changed the carousel filters. Status ok.

4/16/02

MR

The filters looked as if they where not being sampled according to the system e-mails. Changed the filter maintenance switch from run to maintenance for 30 secs then returned it to the run position.

4/12/02

MR

Changed the carousel filters. It looked as if the last 2 filters where not sampled. Status ok.

4/5/02

MR

The filters looked as if they had not been sampled. Returned the filters to the system and flipped the switch to the maintenance position several times and the returned it to the run position.

3/29/02

MR

Changed the carousel filters. Status ok.

3/21/02

 

MR

The system e-mails looked as if the filters where not being sampled. Checked the filters and found that they looked as if they had not been sampled. Flipped the change switch to the change position for 30 secs and then changed it back to the run position. The filters will not be changed this week due to them not being sampled.

3/15/02

MR

Changed the carousel filters. Status ok.

3/8/02

MR

The filters are due to be changed but looked as if the system has not been sampling them. Left the pack in the system and notified the mentor.

3/5/02

MR

The mentor notified me that she was not receiving any volume data. Flipped the change switch for 30 seconds and then returned it to the run position.

3/1/02

MR

Changed the carousel filters. Status ok.

2/22/02

MR

Changed the carousel filters. Status ok.

2/19/02

MR

The mentor notified me that she was not receiving any filter volume data. Flipped the filter change switch for 30 seconds and then returned it to the run position.

2/15/02

MR

Changed the carousel filters. Status ok.

2/15/02

MR

The mentor wants me to check the flow controller. Checked the flow controller and it was reading 19.8 LPM. Notified the mentor and she said it should be reading 30 LPM. Adjusted the flow controller as per instructions received from the mentor back to 30.0 LPM. Notified the mentor.

2/11/02

MR

The power to the site was restored. The system program was started at 1700 GMT.

2/8/02

MR

No filter change due to power outage.

2/1/02

MR

No filter change due to power outage.

1/29/02

MR

The mentor notified me that the flow through the filters is still reading low and wants me to check the impactor to make sure it is assembled correctly. Checked the impactor and it was ok. Notified the mentor.

1/25/02

MR

Changed the carousel filters. Status ok.

1/24/02

MR

Due to not having a spare that will fit the system, the old pump was rebuilt. Installed a rebuild kit in the pump. The vacuum returned to 17 PSI and was steady.

1/23/02

MR

Dick Eagan notified me that the system was reporting 1.3 and 1.4 flow values in the system. Checked the pumps and the main air flow pump had 0.00 vacuum. Due to time restraints, a new pump will be installed on 1/24/02.

1/23/02

MR

The instrument mentor has requested that the pump be checked due to flow in the system has dropped below acceptable values. T/S found that the pump was working but vacuum was low. Further T/S found that the internal carbon vanes were worn. Installed a rebuild kit. The pump checked ok at 17 PSI of vacuum. Notified the mentor. RTS at 1600 GMT.

1/22/02

MR

Dick Eagan notified me that the system did not recognize the filter change on 1/18/02. Flipped the filter maintenance switch and let it sit in the change position for 30 seconds and then flipped it back to run.

1/18/02

MR

Changed the carousel filters. Status ok.

1/11/02

MR

The filters looked as if they had not been sampled. Due to the problem of the impactor not being assembled correctly, The system was reset and filter pack T will be resample.

1/10/02

MR

The instrument mentor reports that the flow rate for the carousel filters is still reading 4.0. Proper reading should be 19.0 or higher. T/S found that the impactor was not properly put back together. Removed the impactor and reassembled correctly. Reinstalled the impactor and the flow returned to 19.0.

1/9/02

MR

The instrument mentor indicates that the flow through the carousel filters is extremely low. T/S found that the vacuum pump is ok. Traced the vacuum line down to the solenoid assembly and disconnected the line at that point. The system flow came back up to 19.8. Reattached the line and the flow returned to 4.8. Flipped the maintenance switch so the bypass solenoid would activate and the flow returned 19.8. The system will reset itself at 0000 GMT. Will observe tomorrow to see if the flow remains at 19.8 or drops low again.

1/4/02

MR

Changed the carousel filters. Status ok.

12/28/01

KT

Changed the carousel filters. The filters change schedule has been switched back to Fridays. This filter pack may have run over by one day.

12/20/01

MR

Changed the carousel filters. Status ok.

12/13/01

MR

Received an e-mail from Dick Eagan stating that the system did not recognize the filter maintenance switch last time the filters where changed and he wanted me to leave the switch in the change position for 30 seconds. I positioned the switch in the change position for 30 seconds then flipped it back. Upon checking the filters it was discovered that the system did not even sample any of the filters. Left the current filters in the system and notified the mentor. Will check on Monday with Dick Eagan to see if the system is going to recognize the filter switch change this time.

12/06/01

MR

Changed the carousel filters. They looked as if they had not been sampled.

11/29/01

MR

Changed the carousel filters.

11/22/01

MR

Due to upcoming holiday season the filter change schedule will be changed to Thursdays. This filter change week may be 1 day short due to this change. Changed the carousel filters. They looked as if they had not been sampled.

11/21/01

MR

Dick Eagan notified me that the software shows the filters have not been sampled at all this week and wants me to cycle the filter change switch. Cycled the change switch as requested at 1510 GMT.

11/16/01

MR

Did not change the filters this week due to Dick Eagan working on the system. He should have the system up by tomorrow and the filters should start to sample.

11/9/01

MR

The Filters looked as if the system stopped sampling. The AOS system is still having software problems and Dick Eagan will be here on 11/10/01 to repair the problem. Will not change the filters this week due to the system being down.

11/5/01

MR

Pat Sheridan notified me that the AOS system had a disk failure over the weekend and that the system is not currently working. This may or may not effect the carousel filters.

11/2/01

MR

Changed out the carousel filters. Filters looked as if they where not sampled.

11/1/01

MR

Dick Eagan notified me that the sampling system does not seem to be running the filter sample program and he is currently working on repairing the problem.

10/26/01

MR

Changed out the carousel filters. Status ok.

10/19/01

MR

Changed out the carousel filters. Status ok.

10/16/01

MR

Richard Eagan requests that the watlow meters be checked due to one of them causing the system to reset periodically. This may or may not cause the filters in the carousel to be reset also.

10/12/01

MR

Due to the system being down the filter change schedule has been changed now to Fridays. Changed out the carousel filters. Status ok.

10/5/01

MR

Pat Sheridan returned the AOS system to normal operating status. Next filter change will be on 10/12/01.

10/2/01

MR

Changed out the carousel filters. Pat Sheridan is on site and has the AOS system shut down performing maintenance.

9/25/01

MR

Changed out the carousel filters. Status ok.

9/18/01

MR

Changed out the carousel filters. Status ok.

9/11/01

MR

Changed out the carousel filters. Status ok.

9/4/01

MR

Changed out the carousel filters. Status ok.

8/28/01

MR

Changed out the carousel filters. Status ok.

8/21/01

MR

Changed out the carousel filters. Status ok.

8/14/01

MR

Changed out the carousel filters. Status ok.

8/7/01

MR

Changed out the carousel filters. The filters looked as if they had not been sampled.

7/31/01

MR

I received an e-mail from Dick Eagan stating that the system is not showing any volume data and wants the flow checked. I checked the flow and it was reading 28.6. Called Dick Eagan and notified him as to what I found. He said he would have to look at the program to see what else could be wrong.

7/31/01

MR

Changed the Carousel fitlers. While changing the filter I noticed that the old fitlers looked as if they where not sampled. Notified the mentor of the problem.

7/23/01

MR

Due to being short handed, the carousel filters where changed out today instead of tomorrow. I will change the filters out on the regular scheduled day on the next filter change. This may allow the filters to run 1 day over their run time.

7/17/01

MR

Changed out the carousel filters. Status O.K.

7/13/01

MR

Checked on the filter change lights and they checked good at this time. RTS at 1430 GMT.

7/12/01

MR

Received an e-mail from Dick Eagan stating he had rebooted the system and wanted me to check the system and confirm whether it was working or not. Checked the system. The maintenance switch seemed to work but the opti 22 indicator lights still where not working. Called Dick Eagan and he said everything seemed to be working ok at this time and the system will began to take samples at 0000 GMT today and that I was to check on the indicator lights in the morning.

7/11/01

MR

Changed out the carousel filters. The system does not seem to be running correctly. Called Dick Eagan and he had me recheck the system fuse and AC voltage and position of the opti 22 switches and they all checked ok. Dick Eagan suspects a software bug and will try and work on it and correct the problem. Will await further instructions from him.

7/9/01

MR

The mentor notified me that the system was not sending the volume information and wants the system checked. Checked the system and found that the opti 22 lights where not lit and when the filter maintenance switch was flipped the indicator light would not come on. Checked the fuse and AC voltage and they checked ok. Notified the mentor and Dick Eagan of the status of the system. Will await further instructions from Dick Eagan.

7/3/01

MR

Changed out the carousel filters. Status O.K.

6/26/01

MR

Changed the carousel filters. Status O.K.

6/19/01

MR

Changed the carousel filters. Status O.K.

6/12/01

MR

Changed the carousel filters. Dates on the filters are correct at this time. RTS at 1400 GMT.

6/11/01

MR

The empty box for the filters has returned. Placed the filters in the box and returned to the mentor. Was not sure of the dates and times these filters needed to be reported as. These filters where believed to have been installed on 5/29/01 and removed on 6/5/01 and where labeled as such.

6/7/01

MR

The mentor reports that the dates are not matching with the samples. This is believed to have been caused by the confusion of sending the empty box back before last filter change. Will make sure the dates are correct at this point on the current filters.

6/5/01

MR

Changed the carousel filters. Found last weeks filter box had been sent in with no filters. Sent an e-mail to the mentor and notified her of the situation. Placed last weeks filters in a plastic bag and left them in the glove box to avoid contamination. Will send them to the mentor when the box returns.

5/29/01

MR

Changed the carousel filters.

5/22/01

MR

Changed the carousel filters.

5/15/01

MR

Changed the carousel filters.

5/8/01

MR

Changed the carousel filters.

5/1/01

MR

Changed the carousel filters. RTS at 1955 GMT.

4/24/01

MR

Performed the carousel filter change. The filter change is back on scheduled for Tuesdays.

4/16/01

MR

Performed the weekly filter change. Changed the filters a day early due to the site being short handed and Ken and Myself being gone all week.

4/10/01

MR

Performed the weekly filter change. Status O.K.

4/3/01

MR

Performed the weekly filter change. RTS at 2115 GMT.

3/27/01

MR

Performed the weekly filter change. RTS at 2115 GMT.

3/20/01

MR

Changed the carousel filters. Status O.K.

3/13/01

MR

Changed the carousel filters. Status O.K. Filter change switch seemed to work OK this week.

3/6/01

MR

Received an email from the mentor requesting that the filters # 4 and 5 where not being sampled and needed to be checked. Since today was the day for filter changing I removed the carousel and observed that the filters looked as if they had been sampled but noticed that the filter change switch was not lit up as usual. Notified the mentor and will await further instruction on the problem. Suspect that the system software may be the problem. Continued the replacement and changing of the filters. RTS the system at 1800 GMT.

2/27/01

MR

Performed the carousel filter change.

2/20/01

MR

Changed out the carousel filters. RTS at 1945 GMT.

2/13/01

MR

Changed the filters. Mentor for the AOS is on site performing maintenance on the AOS system and may have some data interruptions for this system.

2/6/01

MR

Changed the carousel filters. RTS at 1930 GMT.

1/30/01

MR

Performed the carousel filter change. Status O.K. RTS at 2130 GMT.

1/23/01

MR

Performed the carousel filter change. Status O.K.

1/16/01

MR

Performed the carousel filter change. Status O.K. RTS at 2055 GMT.

1/9/01

MR

Performed the carousel filter change. Status O.K. RTS at 2100 GMT.

1/2/01

MR

Performed the carousel filter change. Status O.K. RTS at 2115 GMT.

12/26/00

MR

Performed the carousel filter change. Status O.K. RTS at 2050 GMT. Moved the day of the week this procedure is performed due to holiday schedule. Will be changed out on Tuesdays until further notice.

12/18/00

MR

Performed the carousel filter change. Status O.K. RTS at 2145 GMT.

12/11/00

MR

Performed the carousel filter change. Status O.K. RTS at 2110 GMT.

12/4/00

MR

Performed the carousel filter change. Status O.K. RTS at 2200 GMT.

11/27/00

MR

Performed the carousel filter change. Status O.K. RTS at 2130.

11/20/00

MR

Performed the carousel filter change. Status O.K. RTS at 2205.

11/14/00

MR

Received an e-mail message from the mentor that stated the air flow through the filters was 0. Checked the filters and they seemed to be O.K. Flipped the filter change switch to the change position for 1 min and then flipped it back to the sample position as we had done once before.

11/13/00

MR

Performed the carousel filter change. Status O.K. RTS at 2115.

11/3/00

MR

Performed the carousel filter change. Status O.K. RTS at 2205.

10/27/00

KT

At 1922 GMT flipped the switch to change out the carousal filters. At 1945 GMT change out of filters complete. RTS at 2025 GMT.

10/20/00

MR

Performed the carousel filter change. Status O.K.

10/13/00

MR

Changed out the carousel filters. They all looked as if they where all changed out. RTS at 2110.

10/6/00

MR

Checked the filters and it looked as if the first 2 filters where sampled and the last 2 where sampled only. Went ahead and changed the filters. Dick Eagan notified me as to position the filter change switch in the change position and let it stay for 1 min then switch back to the sample position. Performed this procedure after changing the filters out for this week.

10/3/00

MR

Received an e-mail from the mentor stating that the filters looked as if it had not sampled but 2 filters since last weeks change out. Checked the switch and it checked O.K. Switched it back and forth a couple of times and will await further instructions if this does not work. Will try and locate another switch and change this one out.

9/29/00

MR

Performed the carousel filter change out. RTS at 2005 GMT.

9/22/00

MR

Performed the carousel filter change out. RTS at 2030 GMT.

9/15/00

MR

Performed the carousel filter change out. RTS at 2055 GMT.

9/8/00

MR

The carousel filters looked as if they had been sampled and all was working. Started changing the filters out at 2010 GMT. RTS at 2050 GMT.

9/4/00

MR

Checked the system and the indicator lights were lit and seemed to be working. Will check the filters on Friday to see if they where sampled.

9/1/00

MR

Received an e-mail from Dick Eagan today and said to try and leave the switch in the sample position and when the AOS system reads the switch at 0000 GMT they system should start. Verified that the switch was in the sample position.

8/31/00

MR

The Mentor called and notified me that the filters have not been changing right. Checked the system and all valves where tight and working right. Checked the filters and they where clean and looked unused. The indicator lights that show what filter is working are not lit and look as if the system is dead. Dick Eagan has been notified. The change indicator switch is suspected to be the cause. Flipped the switch back and forth a couple of times and left the old filters in.

8/25/00

MR

Performed the carousel filter change out. Status O.K.

8/18/00

MR

Performed the carousel filter change out. Status O.K.

8/11/00

MR

Performed the carousel filter change out. Status O.K.

8/4/00

MR

Changed out the carousel filters. The inline valve was turned back on. Status O.K.

7/28/00

MR

Due to operator error the filters were not placed back inline last PM. Called the mentor and she said to go ahead and reuse the filters for this weeks change out. Placed the filters back in the system and turned the inline valve back on. RTS at 2045 GMT.

7/21/00

MR

Changed out the carousel filters. RTS at 2035 GMT. Status O.K.

7/14/00

MR

Changed the carousel filters at 2035 GMT. Status O.K.

7/7/00

MR

Changed the carousel filters at 2005 GMT. Status O.K.

6/30/00

MR

Changed the carousel filters at 1945 GMT. Status O.K.

6/23/00

MR

Changed the carousel filters at 1500 GMT. Status O.K.

6/16/00

MR

Changed the carousel filters at 2100 GMT. Status O.K.

6/9/00

MR

Changed the carousel filters at 2010 GMT. Status O.K.

6/2/00

MR

Changed out the carousel filters at 2030 GMT. Status O.K.

5/26/00

MR

Performed the carousel filter change. Status O.K. RTS at 2045 GMT.

5/19/00

MR

Performed the carousel filter change. Status O.K. RTS at 2100 GMT.

5/12/00

MR

Performed the carousel filter change. Status O.K. RTS at 2120 GMT.

5/5/00

MR

Performed the carousel filter change. Status O.K. RTS at 2115 GMT.

4/27/00

MR

Performed the carousel filter change out. Status O.K. RTS at 2055 GMT.

4/20/00

MR

Changed out the carousel filters. Filter #6 looked extremely dirty otherwise status was OK. RTS at 2110 GMT.

4/13/00

MR

Performed the carousel filter change out. Status O.K.

4/6/00

MR

Performed the carousel filter change out. Status O.K.

3/30/00

MR

Performed the filter change out. Several carousel filters where extremely dirty otherwise status was found to be O.K. RTS at 2200.

3/23/00

MR

Performed the filter change out. Status O.K.

3/16/00

MR

Performed the filter change out. Status O.K.

3/9/00

MR

Performed the filter change out. Status O.K.

3/2/00

MR

Changed the carousel filters out. Filter set U was removed and fitler set X was installed. RTS at 2100.

2/24/00

CAM

Pat Sheridan was on site all week and had the system off line much of the time. Pat recommended that the carousel filters not be changed this week.

2/17/00

MR

Changed out the carousel filters. Removed filter set W and installed filter set U. RTS at 2110.

2/10/00

MR

Performed the AOS Carousel filter change.

2/3/00

MR

Performed the AOS Carousel filter change.

1/28/2000

MR

Performed the AOS Carousel filter change. Status O.K. RTS at 2115.

1/20/2000

MR

Dick Eagan and Pat Sheridan were on site the week of 1/10 - 1/14. During that time the AeroChem sampling SW was updated and is now operational. Site Ops will now perform the filter changing procedure on a weekly basis. Filter changing procedure started at 2055 GMT. Completed changes and system RTS at 2200 GMT. Status O.K.

10/22/99

CAM

The Aerosol Chemistry instrument has not yet been fully brought on line for data acquisition. Dick Eagan has some final software preparations to make and Site Operations expect the instrument to be ready for full operation within the next couple of weeks.

7/30/99

MR

The Aerosol Chemistry  instrument has been installed, but no filters have yet been installed.  In support of testing by Dick Eagan for Air Chem data collection, there was a need to install the operational filters in the Air Chem system.

Closed the ball valve to the central line and removed the impactor, then detached the lines into the filter carousel.  Placed the filter carousel and impactor in the glove box.  Using the glove box access, disassembled the carousel and installed the filters 1 through 8.  Placed 2 filters in the impactor.  Reassembled the impactor and reinstalled the impactor and carousel back in the Air Chem equipment rack.  Reattached the lines to the carousel and opened the ball valve in preparation for aerosol sampling.